# Finding Quasars behind the Galactic Plane. IV. Candidate Selection from Chandra with Random Forest

### v1

@@ V.3 Spectroscopic Follow-up of GPQ Candidates
[…]
The NGPS data were reduced with the PypeIt [^3] https://pypeit.readthedocs.io/en/release/index.html spectroscopic reduction pipeline (Prochaska2020), following standard long-slit procedures. Briefly, the two-dimensional frames were bias/overscan corrected and flat-fielded, after which the object trace was determined in each channel. Sky/background was estimated from source-free regions adjacent to the trace and subtracted, and one-dimensional spectra were extracted with a fixed-width aperture centered on the trace. Wavelength solutions were derived from arc-lamp exposures obtained in the standard NGPS calibration sequence. The R- and I-channel spectra were then placed on a consistent relative scale and combined by normalizing in their overlap region to produce a continuous spectrum for each target. Standard-star observations from the same run were used for relative flux calibration and to mitigate telluric absorption features.
@@ VI Summary and Conclusions
We construct a high-quality CSC–WISE–GAIA sample through a series of photometric, astrometric, and morphological quality cuts designed to mitigate source confusion and contamination in crowded fields. Using spectroscopically confirmed quasars and stars compiled from DESI, SDSS, LAMOST, and the literature, we train a RF classifier to distinguish quasars from Galactic sources. To further suppress stellar contamination, we apply a probabilistic zero–proper-motion criterion based on Gaia astrometry. In addition, we estimate photometric redshifts for the quasar candidates using a RF–based regression model incorporating optical, infrared, and X-ray features.
Applying this framework to previously unclassified CSC sources, we identify a total of 6286 quasar candidates across the sky. Among these, 863 sources are located behind the Galactic plane ($|b|<20^{\circ}$) and are identified as GPQ candidates, with 514 classified as high-confidence candidates based on their quasar membership probabilities. Pilot optical spectroscopy confirms two high-confidence GPQ candidates, 2CXO J031404.4+403900 ($z=1.2582$) and 2CXO J031712.7+405301 ($z=1.1313$), providing direct validation of our selection approach. Motivated by these initial confirmations, we will carry out more extensive spectroscopic follow-up of the full GPQ candidate sample.
@@
Comparison with the unified all-sky quasar candidate catalog CatGlobe shows that while the majority of our GPQ candidates have counterparts in existing catalogs, a significant fraction does not. These additional candidates exhibit properties consistent with the broader GPQ population, underscoring the importance of incorporating X-ray information to achieve a more complete census of quasars in the Galactic plane.
@@
Overall, this study demonstrates that X-ray–based selection with Chandra, combined with multi-wavelength data and modern classification techniques, provides an effective pathway for identifying quasars in the most heavily obscured regions of the sky. The resulting GPQ candidate sample represents a valuable resource for future spectroscopic follow-up, for improving the uniformity of all-sky quasar catalogs, and for applications ranging from Galactic absorption studies to the construction of astrometric reference frames in the Galactic disk. Future extensions of this work to larger X-ray datasets and upcoming facilities will further enhance our ability to uncover quasars hidden behind the Milky Way. The significant increase of GPQ sample will help improve the reference frame for astrometry by using GPQs as background references and probe the Milky Way interstellar and circumgalactic media with the absorption features of GPQs.
@@
We acknowledge the supports from National Natural Science Foundation of China under grant Nos. 12133001 and 12573110, as well as from National Key R&D Program of China (2025YFA1614101). We acknowledge the support of Shenzhen Science and Technology program (JCYJ20230807113910021) and the Natural Science Foundation of Top Talent of SZTU (GDRC202208) This research made use of data from the Chandra X-ray Observatory, in particular the Chandra Source Catalog (CSC website) provided by the Chandra X-ray Center (CXC). It also made use of data from the European Space Agency (ESA) mission Gaia (Gaia website), processed by the Gaia Data Processing and Analysis Consortium (DPAC consortium). We used mid-infrared photometry from WISE/NEOWISE via the CatWISE2020 catalog and made use of the NASA/IPAC Infrared Science Archive (IRSA), which is funded by NASA and operated by the California Institute of Technology. Optical photometry was taken from the Pan-STARRS1 Surveys (PS1) and, where PS1 coverage is unavailable, from the NOIRLab Source Catalog DR2 (NSC website). We used dust reddening information based on Planck. We made use of public spectroscopic classifications from DESI DR1, SDSS DR18, and LAMOST; LAMOST is a National Major Scientific Project built by the Chinese Academy of Sciences, funded by the National Development and Reform Commission, and operated and managed by the National Astronomical Observatories, Chinese Academy of Sciences. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France, and services provided by Astro Data Lab, part of the Community Science and Data Center Program of NSF NOIRLab, operated by AURA under a cooperative agreement with the U.S. National Science Foundation. Spectroscopic observations were obtained with the Palomar 200-inch Hale Telescope at Palomar Observatory using the Next Generation Palomar Spectrograph (NGPS). We thank the Palomar Observatory staff for their support during the observations. For DESI and SDSS, we acknowledge the respective collaborations and data release teams, their funding sources, and the observatory sites. In particular, we recognize the cultural significance of I’oligam Du’ag (Kitt Peak) to the Tohono O’odham Nation. This work made use of TOPCAT and scikit-learn for catalog cross-matching and machine-learning analyses, and PypeIt for spectroscopic data reduction.

### v2

@@ V.3 Spectroscopic Follow-up of GPQ Candidates
The most critical next step is spectroscopic confirmation of the candidate GPQs. This will definitively validate the RF method’s accuracy, measure precise redshifts, and allow for detailed studies of the physical properties (e.g., black hole masses, accretion rates) of these previously hidden quasars. To provide an initial validation of our GPQ candidate selection, we obtained pilot optical spectroscopy for two high-confidence GPQ candidates, 2CXO J031404.4+403900 (catalog ) and 2CXO J031712.7+405301 (catalog ). The observations were carried out on 2025 December 14 (UT) with the Palomar 200-inch (P200) Hale telescope using the Next Generation Palomar Spectrograph (NGPS; Jiang et al., 2018). NGPS records simultaneous spectra in two channels, covering approximately 580–780 nm (R channel) and 760–1040 nm (I channel). We used a $1\farcs 5$ slit for both targets.
The NGPS data were reduced with the PypeIt [^2] https://pypeit.readthedocs.io/en/release/index.html spectroscopic reduction pipeline (Prochaska et al., 2020), following standard long-slit procedures. Briefly, the two-dimensional frames were bias/overscan corrected and flat-fielded, after which the object trace was determined in each channel. Sky/background was estimated from source-free regions adjacent to the trace and subtracted, and one-dimensional spectra were extracted with a fixed-width aperture centered on the trace. Wavelength solutions were derived from arc-lamp exposures obtained in the standard NGPS calibration sequence. The R- and I-channel spectra were then placed on a consistent relative scale and combined by normalizing in their overlap region to produce a continuous spectrum for each target. Standard-star observations from the same run were used for relative flux calibration and to mitigate telluric absorption features.
@@ VI Summary and Conclusions
We construct a high-quality CSC–WISE–GAIA sample through a series of photometric, astrometric, and morphological quality cuts designed to mitigate source confusion and contamination in crowded fields. Using spectroscopically confirmed quasars and stellar-type objects compiled from DESI, SDSS, LAMOST, and the literature, we train a RF classifier to distinguish quasars from Galactic sources. To further suppress stellar contamination, we apply a probabilistic zero–proper-motion criterion based on Gaia astrometry. In addition, we estimate photometric redshifts for the quasar candidates using a RF–based regression model incorporating optical, infrared, and X-ray features.
Applying this framework to previously unclassified CSC sources, we identify a total of 7570 quasar candidates across the sky. Among these, 1060 sources are located behind the Galactic plane ($|b|<20^{\circ}$) and are identified as GPQ candidates, with 551 classified as high-confidence candidates based on their quasar membership probabilities ($P_{\mathrm{QSO}}>0.8$). Pilot optical spectroscopy confirms two high-confidence GPQ candidates, 2CXO J031404.4+403900 (catalog ) ($z_{\mathrm{spec}}=1.2582$) and 2CXO J031712.7+405301 (catalog ) ($z_{\mathrm{spec}}=1.1313$), providing direct validation of our selection approach. Motivated by these initial confirmations, we will carry out more extensive spectroscopic follow-up of the full GPQ candidate sample.
@@
Comparison with the unified all-sky quasar candidate catalog CatGlobe shows that while 600 of our GPQ candidates have counterparts in existing catalogs, a substantial fraction (460 sources) are newly identified. These additional candidates exhibit properties consistent with the broader GPQ population, underscoring the importance of incorporating X-ray information to achieve a more complete census of quasars in the Galactic plane.
@@
Overall, this study demonstrates that X-ray–based selection with Chandra, combined with multi-wavelength data and modern classification techniques, provides an effective pathway for identifying quasars in the most heavily obscured regions of the sky. The resulting GPQ candidate sample represents a valuable resource for future spectroscopic follow-up, for improving the uniformity of all-sky quasar catalogs, and for applications ranging from Galactic absorption studies to the construction of astrometric reference frames in the Galactic disk. Future extensions of this work to larger X-ray datasets and upcoming facilities will further enhance our ability to uncover quasars hidden behind the Milky Way. The significant increase of the GPQ sample will help improve the reference frame for astrometry by using GPQs as background references and probe the Milky Way interstellar and circumgalactic media with the absorption features of GPQs.
@@
We are grateful to the reviewer for his valuable comments and constructive suggestions, which have significantly improved the quality of our manuscript. We acknowledge the supports from National Natural Science Foundation of China (NSFC; grant Nos. 12133001, 12573110, 12273076), and National Key R&D Program of China (2025YFA1614101). We acknowledge the support of Shenzhen Science and Technology program (JCYJ20230807113910021) and the Natural Science Foundation of Top Talent of SZTU (GDRC202208) This research made use of data from the Chandra X-ray Observatory, in particular the Chandra Source Catalog (CSC website) provided by the Chandra X-ray Center (CXC). It also made use of data from the European Space Agency (ESA) mission Gaia (Gaia website), processed by the Gaia Data Processing and Analysis Consortium (DPAC consortium). We used mid-infrared photometry from WISE/NEOWISE via the CatWISE2020 catalog and made use of the NASA/IPAC Infrared Science Archive (IRSA), which is funded by NASA and operated by the California Institute of Technology. Optical photometry was taken from the Pan-STARRS1 Surveys (PS1) and, where PS1 coverage is unavailable, from the NOIRLab Source Catalog DR2 (NSC website). We used dust reddening information based on Planck. We made use of public spectroscopic classifications from DESI DR1, SDSS DR18, and LAMOST; LAMOST is a National Major Scientific Project built by the Chinese Academy of Sciences, funded by the National Development and Reform Commission, and operated and managed by the National Astronomical Observatories, Chinese Academy of Sciences. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France, and services provided by Astro Data Lab, part of the Community Science and Data Center Program of NSF NOIRLab, operated by AURA under a cooperative agreement with the U.S. National Science Foundation. Spectroscopic observations were obtained with the Palomar 200-inch Hale Telescope at Palomar Observatory using the Next Generation Palomar Spectrograph (NGPS). We thank the Palomar Observatory staff for their support during the observations. For DESI and SDSS, we acknowledge the respective collaborations and data release teams, their funding sources, and the observatory sites. In particular, we recognize the cultural significance of I’oligam Du’ag (Kitt Peak) to the Tohono O’odham Nation. This work made use of TOPCAT and scikit-learn for catalog cross-matching and machine-learning analyses, and PypeIt for spectroscopic data reduction.